OpenStudy (anonymous):

Alyssa may apply for an unsubsidized Stafford Loan with an interest rate of 6.8%, compounded monthly and a grace period of six months from the time of graduation. Or, Alyssa may have her parents apply for a PLUS Loan with an interest rate of 7.8%, compounded monthly. She will be graduating from college in a year and a half and needs a loan in the amount of $11,895. Which loan is the better choice and why?

OpenStudy (anonymous):

-The Stafford Loan is the better choice because it is $197.81 cheaper than the PLUS Loan. -The PLUS Loan is the better choice because it is $256.32 cheaper than the Stafford Loan. -The PLUS Loan is the better choice because it is $454.13 cheaper than the Stafford Loan. -The Stafford Loan is the better choice because it is $273.50 cheaper than the PLUS Loan.

OpenStudy (anonymous):

@ajprincess @AravindG @Alyx_da_Boss @Anyssah @alahrichi1 @A.kumar

OpenStudy (anonymous):

use ur own brain

OpenStudy (anonymous):

I tried. Give me a formula and I can do it. I can't find the formula.

OpenStudy (mathmate):

Amount after n months, compounded monthly A=P(1+r/12)^n r=interest rate per annum P=principal (borrowed amount) n=number of months money was borrowed. Calculate the amount to be repaid for each option and make your choice accordingly.

OpenStudy (anonymous):

pi=prt/100

OpenStudy (anonymous):

@mathmate what would n be? i put 18, but i don't think that's right.

OpenStudy (mathmate):

I am not too sure if the "grace period" is just an extended repay period where interest continues to grow. If you calculate it for n=18, you would get one of the answers. And for n=24, you would get another answer. The answer depends on when she is ready to repay (completely), 18 or 24 months. Make a choice.

OpenStudy (anonymous):

Grace period is the months where interest isn't charged. Like a credit card.

OpenStudy (anonymous):

I'm not getting the right answer. I'll just pick a random answer.

OpenStudy (mathmate):

If that's the interpretation, then PLUS loan will charge (higher) interest for 24 months while the Stafford loan charges (lower) interest for 18 months. The difference comes to be $727.64, which is not one of the answers.

OpenStudy (anonymous):

The Stafford Loan is the better choice because it is $197.81 cheaper than the PLUS Loan.